Nadyah:

If 3x-y=12, what is the value of 8x divded by 2y

KpopAsianDramaFanYT:

Cancel the common factor of 8 and 2. 4x/y

mypasswordiscool2:

i think 0

mypasswordiscool2:

oh wait

mypasswordiscool2:

nvm u cant say the direct answer

crispyrat:

that is wrong 3x-y=12 so that means 3x=12-y which means x=(12-y)/3 so then (8(12-y)/3)/2y

KpopAsianDramaFanYT:

@crispyrat my answer is not a direct answer

crispyrat:

@kpopasiandramafanyt @mypasswordiscool2 both of you are wrong btw

crispyrat:

@crispyrat wrote:
that is wrong 3x-y=12 so that means 3x=12-y which means x=(12-y)/3 so then (8(12-y)/3)/2y
can you simplfy more?

Nadyah:

i got it thanks

crispyrat:

yw good job!